What is the equation of the line through (8,6) and (2,-4)?
Thanks,

Answer by checkley71(8403)

SLOPE (m)=(Y2-Y1)/(X2-X1)
m=(-4-6)/(2-8)
m=-10/-6
m=10/6
m=5/3 IS THE SLOPE.
NOW REPLACE THE X & Y TERMS IN THE LINE EQUATION WITH ONE SET OF POINTS & SOLVE FOR THE Y INTERCEPT(b).
6=5/3*8+b
6=40/3+b
b=6-40/3
b=(18-40)/3
b=-22/3 ANSWER FOR THE Y INTERCEPT.
THUS THE LINE EQUATION IS:
Y=5X/3-22/3